How does this work?
Choose the ordered pair that is a solution to the system of equations.
-3x + 2y = 10

-2x - y = -5
(-2, 5)
(2, -5)
(0, -5)
(0, 5)

Found 2 solutions by checkley71, fastblue:
Answer by checkley71(8403)

MULTIPLY THE SECOND EQUATION BY 2 & THEN ADD THEM:
-4X-2Y=-10
-3X+2Y=10
---------------------
-7X=0
X=0 ANSWER.
NOW SUBSTITUTE 0 FOR X & SOLVE FOR Y
-3*0+2Y=10
2Y=10
Y=10/2
Y=5 ANSWER.
PROOF
-2*0-5=-5
-5=-5
ANSWER IS (0,5)

Answer by fastblue(13)

-3x + 2y = 10
-2x - y = -5
Multiply 2nd equation by two and add to first
Gives you -5X =0, so x=0, then substitute back in and gives
you y =5, so the ordered pair that solves the equation
is (0,5)
(-2, 5)
(2, -5)
(0, -5)
(0, 5)